OpenStudy (destroyer5):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 3 over 4 power, all over 2 to the 1 over 2 power the eighth root of 2 to the third power - this one! the square root of 2 to the 3 over 4 power the fourth root of 2 the square root of 2

jhonyy9 (jhonyy9):

2^(3/4) -------- = ? 2^(1/2) this is your question ?

OpenStudy (destroyer5):

isnt it number 1?

jhonyy9 (jhonyy9):

so this is equal = 2^(3/4 -1/2) = 2^?

OpenStudy (triciaal):